OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

@geekfromthefutur @kirbykirby @johnweldon1993

OpenStudy (kirbykirby):

Method 1: find the equation of the line yourself: that is, you can find the line in the form y=mx+b easily by finding the slope "m" first... the slope formula being \[ \frac{y_2-y_1}{x_2-x_1}\] where the numerator values are the y-coordinates of your points, and the denominator are the x-coordinates of your points. Once you have "m" you just need to find the y-intercept "b" which you can do by plugging in either of your 2 points into the formula y = mx+b and then solve for "b" Then, you might have to manipulate the equation a bit for it to math the one listed in your choices. Method 2: Just plug in the values of x and y into the line equation for both points. If the equation is satisfied for both points, then both points will be on the line.

OpenStudy (johnweldon1993):

\[\large y - y_0 = m(x - x_0)\] \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

oh ok got it thx guys :D

OpenStudy (kirbykirby):

=]